Nancy is 2 less than twice the age of Sara. The sum of their ages is 28.
A) Write a system of equations that can be used to solve this problem.
B) Solve the system of equations to tell you how old Nancy and Sara each are.

Answer by checkley79(3341)

N=2S-2
N-2S=-2
N+S=28 SUBTRACT.
-----------------------
-3S=-30
S=-30/-3
S=10 ANS. FOR SARA'S AGE.
N=2*10-2
N=20-2=18 AGE OF NANCY.
PROOF:
10+18=28
28=28